A sample of human brain volumes ​(cm3​) is given below. Use the given data values to identify the corresponding z scores that are used for a normal quantile​ plot, then identify the coordinates of each point in the normal quantile plot. Construct the normal quantile​ plot, then determine whether the data appear to be from a population with a normal distribution. 1086 1066 1025 1086 1407 1073 962 1045 List the z scores for the normal quantile plot.
Answer #1

> x=c(1086,1066,1025,1086,1407,1073,962,1045);x
[1] 1086 1066 1025 1086 1407 1073 962 1045
> m=mean(x);m
[1] 1093.75
> s=sqrt(var(x));s
[1] 133.0454
> z=(x-m)/s;z
[1] -0.0582508 -0.2085755 -0.5167410 -0.0582508 2.3544599 -0.1559618 -0.9902637
[8] -0.3664163
> zscores=round(z,2)
> zscores
[1]( -0.06, -0.21, -0.52, -0.06, 2.35, -0.16 ,-0.99, -0.37)
># is the z-score for  normal quantile plot.

# normal quantile plot. is

> qqnorm(x)
>
Normal Q-Q Plot Sample Quantiles 1000 1100 1200 1300 1400 -1.5 -1.0 -0.5 0.0 0.5 1.0 1.5 Theoretical Quantiles